Required information Problem 1-7A Analyzing transactions and preparing financial statements LO C The following information applies to the questions displayed below] Gabi Gram started The Gram Co, a new business that began operations on May 1. The Gram Co. completed the following transactions during its first month of operations. May 1 G.Gram invested $43,000 cash in the company in exchange for its common stock 1 The company rented a furnished office and paid $2,300 cash for Mays rent. 3 The company purchased $5,410 of office equipment on credit. 5 The company paid $730 cash for this months cleaning services. 8 The company provided consulting services for a client and inmediately collected $5,700 cash. 12 The company provided $2,900 of consulting services for a client on credit. 15 The company paid $790 cash for an assistants salary for the first half of this month. 20 The company received $2,900 cash payment for the services provided on May 12. 22 The company provided $3,900 of consulting services on eredit 25 The company received $3,900 cash payment for the services provided on May 22. 26 The company paid $5,410 cash for the office equipment purchased on May 3. 27 The company purehased $85 of advertising in this month (May) local paper on ereditj cash payment ia due June 1 28 The company paid $790 cash for an aseistanto salary for the second halt of this month. 30 The company paid $300 eash for this months telephone bill. 30 The company paid $280 cash for this months uti company paid $1,900 cash in dividends to the ovner (aole shareholder) Problem 1-7A Part 1 Required: 1. Enter the amount of each transaction on individual items of the accounting equation. Do not determine new account balances after each transaction. (Enter the transactions in the given order. Enter reductions to account balances with a minus sign.) Equity Assets Accounts Receivable Liabilities + Accounts CommonDividends Office ent Payable Cash Next > ć Prev 670r7 iii
each transaction. (Enter the transactions in the given order. Enter reductions to account balances with a minus sign.) Assets Equity Office Accounts Date Cash+ May 1 12 15 6 7 of 7 Next> くPrev
Answer #1
Assets = Liabilities + Equity
Date Cash + Accounts
receivable
+ Office
equipment
= Accounts
payable
+ Common
stock
- Dividends + Revenues - Expenses
May 1 43000 + + = + 43000 - + -
1 -2300 + + = + - + - 2300
3 + + 5410 = 5410 + - + -
5 -730 + + = + - + - 730
8 5700 + + = + - + 5700 -
12 + 2900 + = + - + 2900 -
15 -790 + + = + - + - 790
20 2900 + -2900 + = + - + -
22 + 3900 + = + - + 3900 -
25 3900 + -3900 + = + - + -
26 -5410 + + = -5410 + - + -
27 + + = 85 + - + - 85
28 -790 + + = + - + - 790
30 -300 + + = + - + - 300
30 -280 + + = + - + - 280
31 -1900 + + = + - 1900 + -
43000 + 0 + 5410 = 85 + 43000 - 1900 + 12500 - 5275
48410 = 85 + 48325